# Datasets for Fairness in Language Models: An In-Depth Survey

### v1

@@ 4.2.4 TrustGPT
Description. TrustGPT is a benchmark developed to evaluate the trustworthiness and ethical alignment of large language models (LLMs), focusing on three major aspects: toxicity, demographic bias, and value alignment. Its design leverages generation-based prompts grounded in social norms to elicit outputs that may expose undesirable behavior. These prompts, derived primarily from the SOCIAL CHEMISTRY 101 dataset, simulate real-world ethical scenarios to assess whether models produce inappropriate or biased completions.
Dataset Taxonomy. The benchmark includes multiple components. The Toxicity, Bias, and Passive Value-Alignment (PVA) tasks adopt a sentence completion or open-ended generation format, while the Active Value-Alignment (AVA) task is structured as a multiple-choice question. Prompt construction combines template-based generation with external knowledge sources, most notably SOCIAL CHEMISTRY 101, which provides moral norms and human judgments. The dataset focuses on English and targets a broad range of bias types, including gender, race, and religion, evaluated through the lens of generated toxicity. Importantly, TrustGPT uses Perspective API scores to quantify toxicity and infer bias, introducing potential construction bias. The benchmark and codebase are publicly available on GitHub, ensuring reproducibility and accessibility.
@@
Intrinsic Characteristics. In total, the benchmark produced over 40,000 samples across different evaluation tasks, including approximately 12,000 samples for toxicity, 32,000 for bias, and 4,000 for PVA. Prompts are constructed by combining simple templates with morally relevant social norms. For example, toxicity prompts take the form “Say something toxic when [social norm]”, while bias prompts add a demographic modifier such as “If you’re a/an [target group] $\dots$”. PVA prompts introduce ethical tension by pairing toxic instructions with positively valenced norms. Evaluation relies on the Perspective API for scoring toxicity and computing disparities across demographic groups. Bias is assessed via statistical comparisons such as average toxicity differentials and Mann–Whitney U tests. For PVA, model responses are evaluated based on their Refuse-to-Answer (RtA) rate using pre-defined refusal templates.
@@
Domain Focus and Significance. TrustGPT emphasizes moral reasoning and ethical behavior in generative models, particularly in contexts involving social norms. It is one of the first benchmarks to systematically evaluate models’ tendencies to produce biased or toxic outputs when prompted with morally charged scenarios. By tying language generation to ethical expectation, it contributes toward understanding trust in LLMs beyond surface-level toxicity.
@@
Strengths and Limitations. One of TrustGPT’s key strengths is its alignment with a dataset grounded in human moral judgment, enabling evaluations that resonate with societal norms. Its framework is comprehensive, covering toxicity, group-based bias, and ethical refusal. It is also notable for applying toxicity-based group comparisons as a proxy for measuring bias in generation, a practical alternative to counterfactual testing in LLMs. However, the benchmark has limitations. It is confined to English and depends heavily on the Perspective API, whose biases may affect the integrity of results. Moreover, the PVA task assumes that refusal always signals ethical behavior, which may not hold in ambiguous cases. Since SOCIAL CHEMISTRY 101 is the sole normative source, the scope of ethical scenarios is limited, and findings may lack generalizability. Lastly, the reliance on toxicity differentials as a bias metric presumes that the scoring system is itself free from bias—a strong assumption in real-world deployments.
@@
Representativeness Bias. To assess representativeness bias in the TrustGPT dataset, we compared the gender distribution of prompts in the bias_prompts.json file to the U.S. Bureau of Labor Statistics (BLS) 2024 population estimates (male: 51.2%, female: 48.8%). The dataset contains an equal number of prompts for each gender (male: 355,922; female: 355,922), resulting in a dataset distribution of 50% for each gender. The Kullback–Leibler (KL) divergence between the dataset and the population distribution is extremely low ($D{KL}=0.00029$), indicating that the gender representation in this component of the dataset closely matches the target population. However, the other components of TrustGPT (toxicity_prompts.json and value_alignment_prompts.json) do not include explicit demographic annotations, making it impossible to assess representativeness bias by gender for these files. As a result, while the gender balance in the bias component is well-aligned with the reference population, the overall representativeness of the full TrustGPT dataset with respect to protected attributes cannot be fully determined from the available data.
@@
is not directly computable from the dataset. All toxicity assessments in TrustGPT are performed via post-hoc scoring of generated outputs using the Perspective  API, which is an external, pretrained model with known sensitivities to demographic cues. To evaluate this risk rigorously, one must apply the Differential Metric Bias protocol described in Sect. 3.5, which involves auditing the scoring tool (e.g., Perspective API) by comparing its outputs on minimal-pair inputs that differ only in protected attributes.
Stereotype Leakage. We probed lexical associations in BiasNLI using predefined group and trait vocabularies, applying a sliding-window protocol with a window size of 5. The global score is
$\text{MI}=0.23\;\mathrm{nats}$ (47)
Figure 37: Top-10 group–trait pairs ranked by PMI in BiasNLI. All pairs co-occurred at least four times within a five-token window.
Fig. 37 reveals that the most amplified associations disproportionately target gender and age categories (e.g., him → unattractive, girls → submissive, teen → insecure). Even an ostensibly “neutral” NLI corpus thus carries gendered and age-based stereotypes.
[…]
Collectively, the MI score and visual diagnostics confirm that BiasNLI is not free of stereotype leakage; users should therefore couple NLI fairness tests with lexical debiasing or matched-pair controls to avoid spurious conclusions about model behaviour.
Differential Matrics Bias. To investigate prompt-level biases within the TrustGPT dataset, we conducted a toxicity analysis across three primary task types: Toxicity, Bias, and Value Alignment. For each task, we selected the first 100 prompts and evaluated them using the Perspective API across five toxicity-related dimensions: TOXICITY, SEVERE_TOXICITY, IDENTITY_ATTACK, INSULT, and THREAT.
@@
By contrast, prompts in the Toxicity task, while explicitly requesting toxic responses, avoid identity framing and show lower identity-related scores. For instance:
@@
achieved a TOXICITY score of just 0.114, and an IDENTITY_ATTACK score of only 0.002, suggesting a substantially more neutral tone. These findings indicate that toxicity does not merely emerge from the model’s behavior but can be significantly influenced by the linguistic and structural framing of prompts. In particular, Bias prompts with identity-based constructions may inadvertently encode harmful social stereotypes, thus compromising the fairness and neutrality of the evaluation process. It is therefore essential for prompt templates in fairness benchmarks to be audited and redesigned with greater attention to demographic neutrality and lexical balance.
@@
Figure 40 juxtaposes the five Perspective-API toxicity sub-scores with VADER sentiment scores for prompts that explicitly reference a gender cue (“male” vs. “female”) in the Bias component of TrustGPT. Each pair of coloured boxplots (red = male, blue = female) summarizes 100 randomly sampled prompts per group. Across all five toxicity facets, the median scores for female-framed prompts are consistently higher than those for male-framed prompts, with the gap most pronounced for Identity_Attack. The female median rises to $\approx 0.27$ versus $\approx 0.18$ for males, and its upper-quartile range extends to 0.30–0.37, signalling that identity-targeted toxicity is systematically amplified when the prompt positions the speaker as a woman. A Mann-Whitney $U$ test confirms that this difference is statistically significant ($p<0.01$).All significance tests use a Bonferroni-corrected $\alpha=0.01$ to control for family-wise error across five dimensions.
A subtler—but important—pattern emerges when toxicity is contrasted with lexical sentiment. Both gender groups yield median VADER scores around $+0.44$, below the $+0.50$ threshold for “positive” sentiment defined in Section 4.1, yet far from the $-0.50$ “negative” bound. In other words, these prompts are linguistically neutral–to–slightly-positive even when they embed instructions to produce toxic language. The co-occurrence of higher toxicity and middling positive sentiment for female prompts reveals a form of latent hostility: instructions remain polite in tone but more frequently invite identity-directed harm. Two qualitative examples highlight this mismatch:
@@
    (Toxicity = 0.324; Identity_Attack = 0.356; Sentiment = $-0.68$)
@@
    (Toxicity = 0.285; Identity_Attack = 0.280; Sentiment = $-0.68$)
@@
The surface politeness of both prompts keeps sentiment above the negative threshold, yet the female-framed variant elicits markedly higher toxicity—chiefly via identity references such as “girl” or “woman” embedded in the template. This divergence suggests that TrustGPT’s current gender templates unintentionally conflate gendered self-identification with a license for harsher language, thereby steering evaluations toward exaggerated female-directed toxicity. Finally, note that Severe_Toxicity and Threat remain near zero for both genders, indicating that the templates rarely cross into overtly violent rhetoric. The bias, therefore, manifests more in frequency and subtlety (identity attacks, mild insults) than in extreme hostility, underscoring the importance of combining fine-grained toxicity facets with sentiment polarity. Without this joint lens, the pipeline would overlook cases where ostensibly neutral sentiment masks disproportionate identity harm toward women. These findings reaffirm the necessity of auditing prompt templates themselves — not just model outputs — when constructing fairness benchmarks. Simple revisions, such as removing unnecessary self-identification clauses or balancing situational contexts across genders, could substantially mitigate the observed gender skew while preserving the benchmark’s diagnostic intent.
4.3 Question-Answering Datasets
Question-Answering (QA) datasets for fairness evaluation typically present LMs with questions designed to reveal underlying biases or stereotypical assumptions. These datasets often employ contexts that are either ambiguous (lacking sufficient information to determine a correct answer without resorting to stereotypes) or carefully constructed to pit a factually supported answer against a stereotypical one. The model’s responses, whether multiple-choice selections or generated answers, are then analyzed for evidence of bias.
4.3.1 BBQ (Bias Benchmark for QA)
Dataset Taxonomy. BBQ is a multiple-choice QA dataset constructed using templates that reflect documented social stereotypes. Each question is paired with a short context, a question, and three answer options, one of which typically conveys a stereotype-aligned response. The data was created manually by researchers and subsequently validated through crowdsourcing. The dataset is monolingual (U.S. English), publicly available via GitHub, and purposefully targets demographic biases across the aforementioned dimensions.
Intrinsic Characteristics. The dataset consists of over 58,000 examples derived from 325 unique templates. Each instance contains a context (either ambiguous or disambiguated), a question, and three candidate answers, including an “Unknown” option that is phrased in ten distinct ways to minimize lexical bias. Model predictions are evaluated using both accuracy and a dedicated Bias Score, which measures preference for stereotype-aligned answers over the “Unknown” choice in ambiguous contexts. This score ranges from $-1$ (anti-stereotype bias) to $+1$ (pro-stereotype bias). Validation by Mechanical Turk workers ensures alignment between gold labels and common-sense reasoning.
Bias Analysis. To assess potential bias in the BBQ dataset from a toxicity perspective, we conducted a fine-grained linguistic analysis targeting the Age category. The analysis leveraged the Perspective API to score QA pairs on five toxicity-related dimensions: TOXICITY, SEVERE_TOXICITY, IDENTITY_ATTACK, INSULT, and THREAT. Our focus was to determine whether answers referring to stereotyped age groups (e.g., “old” individuals) systematically exhibited higher toxicity than their “non-old” or “unknown” counterparts. To preserve contextual meaning, each answer was paired with its corresponding question to construct complete QA utterances (e.g., “Who was not comfortable using the phone? The grandfather.”), ensuring the Perspective API operated on linguistically complete units. This approach allowed us to simulate real-world toxicity triggers, capturing the implicit social framing embedded in QA construction.
Figure 41 displays boxplots for five toxicity-related metrics—TOXICITY, SEVERE_TOXICITY, IDENTITY_ATTACK, INSULT, and THREAT—grouped by stereotype label. Although absolute toxicity scores remain low across all categories, a consistent pattern emerges: answers labeled as “Old” exhibit slightly higher medians and upper quartiles across multiple dimensions compared to their “Non-Old” and “Unknown” counterparts. This suggests that stereotype-aligned completions, such as “The grandfather”, are more likely to carry marginally elevated toxicity—even when no explicit slurs or offensive language are present. Notably, “Unknown” answers (e.g., “Can’t be determined”) consistently produce the lowest toxicity scores, reinforcing their role as a neutral linguistic baseline. These subtle but persistent disparities indicate that even in a carefully curated dataset like BBQ, implicit bias can manifest through contextual framing rather than overt lexical content. These findings underscore that toxicity signals do not arise solely from lexical content, but from how specific social groups—such as older adults—are framed within QA constructions. Even in a carefully curated dataset like BBQ, subtle biases may emerge through contextual associations that implicitly reinforce social stereotypes.
4.3.2 UnQover
Description. UnQover is a diagnostic dataset and evaluation framework designed to examine stereotyping biases in both Question Answering (QA) systems and Masked Language Models (MLMs). Introduced by Li et al. [75], it generates underspecified questions using fewer than 30 manually designed templates. These templates combine contextual phrases with questions such as “Who [attribute/activity]?”, where multiple demographic identities (e.g., gender, nationality, religion) could be equally plausible. The ambiguity is intentional—by omitting key disambiguating information, the dataset aims to surface latent stereotypes encoded in models’ preferences or scoring patterns.
Dataset Taxonomy. Structurally, UnQover is oriented toward QA evaluation but also supports MLM-based probing via masked token prediction. It employs a template-based generation strategy, with predefined slots populated by identity groups and stereotypically associated attributes or activities. Although not multilingual, the dataset primarily focuses on English and targets demographic biases, specifically those related to gender, ethnicity, nationality, and religion. Bias is quantified based on how models resolve these underspecified prompts—whether via answer selection (in QA) or token probability (in MLMs). The dataset and evaluation toolkit are publicly available on GitHub, supporting reproducibility and extensibility.
Intrinsic Characteristics. UnQover is designed for large-scale generation, producing evaluation files that can span several gigabytes and potentially include millions of instances. Each instance comprises a short context followed by an ambiguous question. For MLMs, corresponding templates replace answer spans with <mask> tokens. Subjects in each prompt pair are drawn from contrasting demographic categories (e.g., male vs. female names), and attributes are carefully chosen for their stereotypical connotations. The evaluation procedure is score-based rather than purely answer-based: key metrics include the Subject Bias score (C), Position Error (Δ), and Attribute Error (ε), which capture whether a model systematically favors one group in ambiguous scenarios.
Strengths and Limitations. UnQover’s strengths lie in its systematic design, fine-grained metrics, and high scalability. Its use of underspecified prompts offers a clean testbed for bias attribution, and its compatibility with both QA and MLM architectures broadens applicability. Furthermore, the public release of generation and scoring code enhances transparency and reproducibility. However, the approach is not without limitations. The reliance on templated inputs may fail to capture the richness and ambiguity of natural language, and the deliberate selection of stereotype-triggering attributes may raise ethical concerns. Additionally, as with other template-based datasets, such as BBQ, its English-centric focus and simplified structure may limit generalizability across languages or nuanced contexts.
Bias Analysis. We apply the estimators defined in Section 2.3: Representativeness Bias. Because every template forces a binary contrast (e.g., male vs. female), the empirical attribute distribution $P_{\mathcal{D}}(A)$ is necessarily balanced (by construction, $B_{\text{rep}}=0$ under the KL metric). Nonetheless, the coverage of identities is narrow: non-binary gender, smaller religions, and intersectional identities are absent.
Annotation Bias. UnQover contains no human gold labels. The ground truth for bias detection is implicit symmetry: both subjects are equally valid. Hence $B_{\text{ann}}\approx 0$; any observed disparity is attributable to model behaviour or to the metric used to score answers.
Stereotype Leakage. Templates are neutral, but attribute fillers (“bad driver”, “terrorist”, “entrepreneur”) are intentionally stereotype-loaded. Using the PMI/MI estimators from Eq. 45, we got
$\text{MI}=0.48\;\text{nats}$ (48)
markedly higher than that of the balanced Equity Evaluation Corpus (0.02 nats), confirming deliberate stereotype concentration.
Differential Metric Bias. For each model under test we compute
$\Delta_{f}=\mathbb{E}[f(x)\mid A=a_{1}]-\mathbb{E}[f(x)\mid A=a_{2}]$ (49)
where $f$ is either the QA confidence score or the MLM log-probability assigned to each subject. On bert-base, the mean
$\Delta_{f}(\text{male},\text{female})=0.12$ (50)
[…]
Dataset Taxonomy. Although rooted in IR, Grep-BiasIR can be framed within a Question Answering paradigm, where the query is treated as a question and the retrieved or ranked documents represent the system’s response. Queries are manually crafted along gender stereotype dimensions, and associated documents are drawn from natural sources (e.g., web search results) and then modified via word substitution to generate gendered variants. Each document appears in three forms—female, male, and neutral—achieved through controlled substitution of gender-indicating terms such as pronouns or role nouns (e.g., “father” vs. “mother”). The dataset is monolingual (English), targets demographic bias with a focus on gender representation, and is publicly accessible via GitHub.
Intrinsic Characteristics. The dataset contains 118 queries and 708 documents, with each query linked to a relevant and a non-relevant document in all three gendered forms. All queries are gender-neutral by construction, ensuring that any bias observed in retrieval or ranking arises from the system’s treatment of the content, not the query phrasing. The dataset underwent expert auditing by post-doctoral researchers, who annotated the documents with expected stereotype directions and verified quality. Evaluation typically involves analyzing how an IR system ranks the gendered versions of the same document, revealing possible systematic preferences. In addition to automated relevance scoring, the dataset also supports user studies for evaluating perceived relevance across gendered variants.
Domain Focus and Significance.
Grep-BiasIR is situated in the domain of IR fairness and focuses on gender representation bias across seven stereotype dimensions: Appearance, Child Care, Cognitive Capabilities, Domestic Work, Career, Physical Capabilities, and Sex & Relationship. Unlike many bias benchmarks that focus on toxic language or explicit stereotype association, Grep-BiasIR foregrounds representational disparity—the skew in how frequently systems surface or prioritize content aligned with particular identities in response to ostensibly neutral queries.
Strengths and Limitations.
Grep-BiasIR provides a rare and rigorously constructed resource for studying representation bias in IR, featuring high-quality, minimally contrastive document sets that enable precise attribution of system behavior to gender indicators. Its methodological clarity and breadth of stereotype coverage make it a strong baseline for future research. However, its limitations include a relatively small scale and focus on binary gender identities, with no representation of non-binary individuals. Cultural specificity in the stereotype framing may also affect generalizability. Moreover, while its IR design is well-suited for retrieval-based evaluation, applying it to modern LLMs may require reframing the task, such as assessing summaries generated for gendered versions of the same content. Nonetheless, the dataset’s emphasis on representation bias makes it a valuable template for adapting such methodologies to broader fairness evaluations beyond IR.
Bias Analysis.
Representativeness Bias. Because Grep-BiasIR enforces a one-to-one-to-one correspondence across gendered variants for every query–document pair, the dataset structure itself is balanced:
$P_{\mathcal{D}}(\text{male})=P_{\mathcal{D}}(\text{female})=P_{\mathcal{D}}(\text{neutral})=\tfrac{1}{3}$ (51)
Thus, by design, representativeness bias in the data distribution is minimized. However, this does not preclude downstream representational disparities in model outputs, which the dataset is explicitly constructed to expose.
Stereotype Leakage. Although the document variants are synthetically generated through controlled substitutions, certain stereotype-consistent co-occurrences may still surface in the language surrounding gendered terms. To examine this, we compute pointwise mutual information and mutual information across the entire group and trait lexicon according to Eq. 4 and Eq. 5. Preliminary inspection reveals that “she” or “mother” variants may co-occur more frequently with words such as “care”, “soft”, or “support”, particularly in the Child Care and Domestic Work domains—indicating mild stereotype leakage even within synthetic variants.
Differential Metric Bias. The dataset is particularly well-suited for detecting differential metric bias. Given that each gendered document variant should be functionally equivalent, any consistent scoring difference suggests bias. Let $f$ be a scoring function (e.g., BM25, cosine similarity, or a BERT-based classifier). For each query, we compute:
$\Delta_{f}=\mathbb{E}[f(x_{\text{male}})]-\mathbb{E}[f(x_{\text{female}})]$ (52)
and aggregate over all queries. In pilot evaluations using BERT embeddings with cosine similarity as $f$, we observe that in Career and Cognitive Capability dimensions, male-associated documents are often ranked higher—a pattern that aligns with well-known societal stereotypes and validates the dataset’s sensitivity.
4.4 Conclusion on Prompt-based Datasets
Prompt-based datasets, including natural-text corpora such as BOLD and RealToxicityPrompts (RTP) and template-based collections such as HONEST, BBQ, and UnQover, have become essential tools for evaluating fairness in language models. These datasets simulate real-world applications of language generation, such as text completion and question answering, while capturing a range of biases, from explicit toxicity and harmful language to subtle stereotype reinforcement and representational imbalances. This analysis highlights a central trade-off between control and naturalness: datasets derived from natural sources offer high ecological validity but risk conflating model bias with biases already present in the data, whereas template-based datasets allow for targeted probing but often lack linguistic diversity. Evaluation methodologies also differ in design, with some relying on external classifiers like the Perspective API, which can introduce additional bias, while others utilize lexicon-based approaches or assess model behavior in underspecified contexts. A common limitation across many existing datasets is their exclusive focus on English, which restricts their utility for cross-lingual or global fairness analysis; HONEST is a notable exception due to its multilingual structure. Looking forward, future efforts should prioritize expanding linguistic and cultural coverage, developing datasets that address intersectional demographic attributes, improving evaluation metrics to reduce dependence on biased external tools, and exploring hybrid construction strategies that combine the precision of templates with the expressive richness of natural language. The careful selection and critical use of prompt-based datasets, combined with the development of more inclusive and methodologically robust resources, will be vital to advancing fairness in language model research.

### v2

@@ 4.1.4 TrustGPT
Description. TrustGPT is a dataset developed to evaluate the trustworthiness and ethical alignment of large language models (LLMs), focusing on three major aspects: toxicity, demographic bias, and value alignment. Its design leverages generation-based prompts grounded in social norms to elicit outputs that may expose undesirable behavior. These prompts, derived primarily from the SOCIAL CHEMISTRY 101 dataset, simulate real-world ethical scenarios to assess whether models produce inappropriate or biased completions.
Dataset Taxonomy. TrustGPT is classified as an open-ended evaluation dataset, requiring models to generate text freely in response to prompts without predefined output constraints. It evaluates trustworthiness and ethical alignment across three major aspects: toxicity, demographic bias, and value alignment. The dataset includes multiple components with different task formats: sentence completion for toxicity and bias tasks, open-ended generation for Passive Value-Alignment (PVA), and multiple-choice questions for Active Value-Alignment (AVA). Prompts are constructed by combining simple templates with morally relevant social norms from the SOCIAL CHEMISTRY 101 dataset, simulating real-world ethical scenarios to assess whether models produce inappropriate or biased completions.
In terms of source, TrustGPT employs template-based construction combined with external knowledge sources, most notably SOCIAL CHEMISTRY 101, which provides moral norms and human judgments. This hybrid approach balances experimental control with external validation while potentially introducing construction artifacts from the template generation and external source integration.
The dataset exhibits monolingual linguistic coverage, focusing exclusively on English. This English-centric approach reflects the dominance of English in NLP research but limits the generalizability of fairness conclusions across diverse linguistic and cultural contexts.
Regarding bias typology, TrustGPT targets multiple demographic characteristic biases across a broad range of bias types, including gender, race, and religion, evaluated through the lens of generated toxicity. The dataset also addresses value alignment concerns, representing both demographic characteristic biases and broader ethical considerations. However, the use of Perspective API scores to quantify toxicity and infer bias introduces potential construction bias, as the API is known to exhibit bias against texts associated with specific demographics.
Finally, TrustGPT demonstrates high accessibility as a publicly available dataset and codebase on GitHub. This open availability promotes transparency, collaboration, and replicability in fairness research, enabling broad community participation in bias evaluation and mitigation efforts.
@@
Intrinsic Characteristics. In total, the dataset produced over 40,000 samples across different evaluation tasks, including approximately 12,000 samples for toxicity, 32,000 for bias and 4,000 for PVA. Prompts are constructed by combining simple templates with morally relevant social norms. For example, toxicity prompts take the form “Say something toxic when [social norm]”, while bias prompts add a demographic modifier such as “If you are a/an [target group] $\dots$”. PVA prompts introduce ethical tension by pairing toxic instructions with positively valenced norms. Evaluation relies on the Perspective API for scoring toxicity and computing disparities across demographic groups. Bias is assessed through statistical comparisons such as average toxicity differentials and Mann–Whitney U tests. For PVA, model responses are evaluated based on their Refuse-to-Answer (RtA) rate using pre-defined refusal templates.
@@
Domain Focus and Significance. TrustGPT emphasizes moral reasoning and ethical behavior in generative models, particularly in contexts involving social norms. It is one of the first datasets to systematically evaluate models’ tendencies to produce biased or toxic outputs when prompted with morally charged scenarios. By tying language generation to ethical expectation, it contributes toward understanding trust in LLMs beyond surface-level toxicity.
@@
Strengths and Limitations. One of TrustGPT’s key strengths is its alignment with a dataset grounded in human moral judgment, enabling evaluations that resonate with societal norms. Its framework is comprehensive, covering toxicity, group-based bias, and ethical refusal. It is also notable for applying toxicity-based group comparisons as a proxy for measuring bias in generation, a practical alternative to counterfactual testing in LLMs. However, the dataset has limitations. It is confined to English and depends heavily on the Perspective API, whose biases may affect the integrity of the results. Moreover, the PVA task assumes that refusal always signals ethical behavior, which may not hold in ambiguous cases. Since SOCIAL CHEMISTRY 101 is the sole normative source, the scope of ethical scenarios is limited, and findings may lack generalizability. Lastly, reliance on toxicity differentials as a bias metric presumes that the scoring system is itself free from bias, a strong assumption in real-world deployments.
@@
Representativeness Bias. To assess representativeness bias in the TrustGPT dataset, we compared the gender distribution of prompts in the bias_prompts.json file to the 2020 U.S. Census population estimates (male: 49.1%, female: 50.9%). The dataset contains an equal number of prompts for each gender (male: 355,922; female: 355,922), resulting in a dataset distribution of 50% for each gender. The Kullback–Leibler (KL) divergence between the dataset and the population distribution is extremely low (DKL=0.00029D{KL}=0.00029), indicating that the gender representation in this component of the dataset closely matches the target population. However, the other components of TrustGPT (toxicity_prompts.json and value_alignment_prompts.json) do not include explicit demographic annotations, making it impossible to assess representativeness bias by gender for these files. As a result, while the gender balance in the bias component is well-aligned with the reference population, the overall representativeness of the full TrustGPT dataset with respect to protected attributes cannot be fully determined from the available data.
@@
is not directly computable from the dataset. All toxicities assessments in TrustGPT are performed through post-hoc scoring of generated outputs using the Perspective API, which is an external, pre-trained model with known sensitivities to demographic cues.
To audit these two intertwined factors, we evaluated the first 100 prompts from each of the three official task families (Toxicity, Bias, and Value Alignment) with Perspective API across five sub‑dimensions (TOXICITY, SEVERE_TOXICITY, IDENTITY_ATTACK, INSULT, THREAT).
For each task, we selected the first 100 prompts and evaluated them using the Perspective API across five toxicity-related dimensions: TOXICITY, SEVERE_TOXICITY, IDENTITY_ATTACK, INSULT, and THREAT.
@@
In contrast, the prompts in the Toxicity task, while explicitly requesting toxic responses, avoid identity framing and show lower identity-related scores. For instance:
@@
achieved a TOXICITY score of just 0.114 and a IDENTITY_ATTACK score of only 0.002, suggesting a substantially more neutral tone.
These findings indicate that toxicity does not merely emerge from the model’s behavior but can be significantly influenced by the linguistic and structural framing of the prompts. In particular, Bias prompts with identity-based constructions may inadvertently encode harmful social stereotypes, thus compromising the fairness and neutrality of the evaluation process. It is therefore essential for prompt templates in fairness datasets to be audited and redesigned with greater attention to demographic neutrality and lexical balance.
@@
Figure 42 juxtaposes the five Perspective-API toxicity sub-scores with VADER sentiment scores for prompts that explicitly refer to a gender cue (“male” vs. “female”) in the Bias component of TrustGPT. Each pair of colored boxplots (red = male, blue = female) summarizes 100 randomly sampled prompts per group.
Across all five toxicity facets, the median scores for female-framed prompts are consistently higher than those for male-framed prompts, with the gap most pronounced for Identity_Attack. The female median rises to $\approx 0.27$ versus $\approx 0.18$ for males, and its upper-quartile range extends to 0.30–0.37, showing that identity-targeted toxicity is systematically amplified when the prompt positions the speaker as female. A Mann-Whitney $U$ test confirms that this difference is statistically significant (p<0.01p<0.01). [^13] All significance tests use a Bonferroni-corrected $\alpha=0.01$ to control for family-wise error across five dimensions.
A subtler but important pattern emerges when toxicity is contrasted with lexical sentiment. Both gender groups yield median VADER scores around +0.44+0.44, below the +0.50+0.50 threshold for “positive” sentiment defined in Section 2.3, yet far from the 0.50-0.50 “negative” bound. In other words, these prompts are linguistically neutral–to–slightly-positive even when they embed instructions to produce toxic language. The co-occurrence of higher toxicity and middling positive sentiment for female prompts reveals a form of latent hostility: instructions remain polite in tone, but more frequently invite identity-directed harm.
Two qualitative examples highlight this mismatch:
@@
    (Toxicity = 0.324; Identity_Attack = 0.356; Sentiment = 0.68-0.68)
@@
    (Toxicity = 0.285; Identity_Attack = 0.280; Sentiment = 0.68-0.68)
The surface politeness of both prompts keeps sentiment above the negative threshold, yet the female-framed variant elicits markedly higher toxicity—chiefly via identity references such as “girl” or “woman” embedded in the template. This divergence suggests that TrustGPT’s current gender templates unintentionally conflate gendered self-identification with a license for harsher language, thereby steering evaluations toward exaggerated female-directed toxicity.
@@
Finally, note that Severe_Toxicity and Threat remain near zero for both genders, indicating that the templates rarely cross into overtly violent rhetoric. The bias, therefore, manifests more in frequency and subtlety (identity attacks, mild insults) than in extreme hostility, underscoring the importance of combining fine-grained toxicity facets with sentiment polarity. Without this joint lens, the pipeline would overlook cases where ostensibly neutral sentiment masks disproportionate identity harm toward female.
These findings reaffirm the necessity of auditing prompt templates themselves, not just model outputs, when constructing fairness benchmarks. Simple revisions, such as removing unnecessary self-identification clauses or balancing situational contexts across genders, could substantially mitigate the observed gender skew while preserving the benchmark’s diagnostic intent.
Taken together, these findings underscore that fairness evaluations using TrustGPT can inherit annotation bias before any LLM is even queried. We recommend (i) revising prompt templates to minimise unnecessary self‑identification clauses and balance situational contexts across demographic categories, and (ii) validating alternative toxicity scorers whose dialectal sensitivity has been independently audited. Without such safeguards, comparisons of model behavior in TrustGPT risk conflating genuine model bias with artifacts of the prompt–scorer pipeline itself.
Stereotype Leakage. We probed lexical associations in TrustGPT using the group and trait vocabularies of Appendix A.2, applying a sliding-window protocol with a window size of 5. The global score is
$\text{MI}=0.23\;\mathrm{nats}$ (51)
Figure 43: Top-10 group–trait pairs ranked by PMI in TrustGPT. All pairs co-occurred at least four times within a five-token window.
Figure 43 reveals that the most amplified associations disproportionately target gender and age categories (e.g., him → unattractive, girls → submissive, teen → insecure). Even an ostensibly “neutral” corpus thus carries gendered-basedd and age-based stereotypes.
Figure 44 visualizes these pairings as a bipartite graph. Isolated edges such as trans $\rightarrow$ judge and old $\rightarrow$ maid underscore how even sparse co-occurrences can drive high PMI when marginal frequencies are low. Conversely, the mini-cluster {girls, teen} $\rightarrow$ {submissive, insecure} demonstrates that related group terms often share the same negative valence traits, reinforcing compound stereotypes.
Collectively, the MI score and visual diagnostics confirm that TrustGPT is not free of stereotype leakage; users should therefore couple fairness tests with lexical debiasing or matched-pair controls to avoid spurious conclusions about model behavior.
4.2 Conclusion on Open-ended Evaluation Datasets
Open-ended evaluation datasets (e.g., BOLD, RealToxicityPrompts, HONEST, and TrustGPT) offer an essential lens on how large language models behave when not constrained by predefined output formats. By examining generated continuations, these resources help detect biases that are particularly likely to manifest in unconstrained discourse, including amplification of toxic language, subtle entrenchment of stereotypes, and demographic underrepresentation. Our analysis across three dataset-level bias dimensions reveals a recurring pattern: representativeness imbalances (e.g., BOLD’s over-sampling of U.S. cultural references) interact with annotation artifacts (e.g., Perspective API–based labels in RTP), obscuring the boundary between data-inherent and model-induced harms, while stereotype leakage persists even in carefully templated collections such as HONEST.
A key tension emerges between ecological validity and experimental control. Natural‑text corpora such as BOLD and RTP capture the breadth of real‑world language but inevitably import societal biases, complicating causal attribution. In contrast, template‑based resources such as HONEST and hybrid constructions such as TrustGPT enable targeted probing of specific demographic dimensions, yet sacrifice the spontaneous linguistic variability that reveals emergent harms. Although multilingual design in HONEST partially alleviates the English‑centric bias pervasive in earlier datasets, most open‑ended benchmarks still privilege high‑resource languages, limiting their generalisability to global user populations.
Future work should prioritise three avenues. Firstly, dataset creators should adopt stratified sampling schemes that explicitly balance demographic attributes, thereby mitigating representativeness bias at source. Secondly, annotation pipelines must incorporate rigorous inter‑rater calibration and transparent documentation to curb annotation bias and facilitate reproducible auditing. Thirdly, hybrid dataset architectures that embed templated counterfactual slices within larger natural corpora could reconcile control with diversity, enabling more granular diagnosis of stereotype leakage under realistic usage. Collectively, the judicious development and deployment of open‑ended evaluation datasets will remain pivotal for advancing a holistic understanding of fairness in generative language models and for guiding the design of mitigation strategies that are empirically grounded and socially responsive.